OpenStudy (mathmath333):

find the sum of the series \(\large\tt \color{black}{1.2+2.2^2+3.2^3....100.2^{100}}\)

OpenStudy (mathmath333):

\(\Large\tt \color{black}{\sum\limits_{n=1}^{100}n.2^n}\) = ?

OpenStudy (freckles):

that is a huge sum

OpenStudy (mathmath333):

yes but there must be formula

OpenStudy (freckles):

i will try to think on it

OpenStudy (mathmath333):

@kirbykirby

OpenStudy (freckles):

no luck

OpenStudy (mathmath333):

\(\Large\tt \color{black}{\sum\limits_{n=1}^{100}n.2^n=(n-1)\times2^{101}+2}\) see if it is correct

OpenStudy (mathmath333):

\(\Large\tt \color{black}{\sum\limits_{n=1}^{100}n.2^n=(99)\times2^{101}+2}\)

OpenStudy (freckles):

Proof by Induction: For n=1 we have 1*2^1=2=(1-1)*2^(1+1)+2 true Now assume for some integer k we have \[\sum_{i=1}^{k}i 2^i=(k-1)2^{k+1}+2\] \[\sum_{i=1}^{k+1}i2^i=(k+1)2^{k+1}+\sum_{i=1}^{k}i2^i \\= (k+1)2^{k+1}+(k-1)2^{k+1}+2 \\=2^{k+1}(k+1+k-1)+2\\=2^{k+1}(2k)+2 \\= k2^{k+2}+2\] There for all integer n we have \[\sum_{i=1}^{n}i2^i=(n-1)2^{n+1}+2\]

OpenStudy (freckles):

how did you come up with that

OpenStudy (mathmath333):

i used the options given ,but i still cant get it a.) \(\Large\tt \color{black}{=(100)\times2^{101}+2}\) b.) \(\Large\tt \color{black}{=(99)\times2^{100}+2}\) c.) \(\Large\tt \color{black}{=(99)\times2^{101}+2}\) d.) none

OpenStudy (freckles):

\[1\cdot2+2\cdot 2^2+3\cdot 2^3\] say we wanted to look at this instead... \[1\cdot 2+2\cdot 2^2+3 \cdot 2^3 \\ \\ \text{ I wonder if we can somehow } \\ \text { consider the geometric sequences below \to help} \\ 1 \cdot 2 +1 \cdot 2^2+ 1 \cdot 2^3=\frac{1-2^{3+1}}{1-2}=1(2^{3+1}-1) \\ 2 \cdot 2+ 2 \cdot 2^2+ 2 \cdot 2^3=2 \frac{1-2^{3+1}}{1-2}=2( 2^{3+1}-1) \\ 3 \cdot 2 +3 \cdot 2^2+3 \cdot 2^3=3\frac{1-2^{3+1}}{1-2}=3(2^{3+1}-1)\]

OpenStudy (freckles):

we have our terms that we care about occurring on that diagonal there

OpenStudy (freckles):

maybe we need to find some pattern for the numbers off the diagonal

OpenStudy (freckles):

so we can subtract them out easily

OpenStudy (mathmath333):

yes u r right

OpenStudy (mathmath333):

but is it possible

OpenStudy (freckles):

just thinking and writing my thoughts here r1+2^4=r2

OpenStudy (freckles):

so maybe we don't need row 2

OpenStudy (freckles):

\[3(2^{3+1}+1)-1(2^{3+1}+1)\]

OpenStudy (freckles):

that should be equal to \[(3-1)2^{3+1}+2\]

OpenStudy (freckles):

ok yeah this is it

OpenStudy (freckles):

nope got lost again

OpenStudy (mathmath333):

but how can u say that we dont need second row

OpenStudy (freckles):

because we can extend the finite series in the first row to get all the terms in the second row

OpenStudy (freckles):

I will continue to think on this i have to go and that is just a whole bunch of me thinking this problem is probably way easier than i'm making it

OpenStudy (mathmath333):

ok thanks very much

OpenStudy (freckles):

best I can come with his observing a pattern in the smaller series and formulate the conjecture (you can also choose to prove your conjecture if you are that level) it or even easier just trying your little formulas to see which works sorry i couldn't be more use

OpenStudy (freckles):

omg that was way easier than anything i was thinking

OpenStudy (mathmath333):

hey but i didnt get that one can u explain that with the link

OpenStudy (mathmath333):

S = -(1.2 + 1.2^2 + ..............+2^100) + 100.2^101 ???

OpenStudy (mathmath333):

ohh ok that leaves(2-1) in bracket

OpenStudy (mathmath333):

u know in exams people get hardly 2 min per this types of questions with options

OpenStudy (mathmath333):

hey can that logic be used to derive formula too

OpenStudy (freckles):

pretend we have the little series we were talking about earlier \[\text{ Let } S=1 \cdot 2 +2 \cdot 2^2+3\cdot 2^3 \\ \text{ then } 2S=1 \cdot (2 \cdot 2 )+2 \cdot (2^2 \cdot 2)+3 \cdot (2^3 \cdot 2) \\ \text{ so subtracting S from 2S will give us S, the \sum we are looking for} \\ 2S-S=1(2^2)+2(2^3)+3(2^4)-[1(2)+2(2^2)+3(2^3)] \\2S-S=-1(2)+(1-2)2^2+(2-3)2^3+3(2^4)\]

OpenStudy (freckles):

like they just put the "like terms together" the parts that had the same exponent

OpenStudy (freckles):

\[2S-S=-1(2)+(-1)2^2+(-1)2^3+3(2^4) \\2S-S=(-1)[2+2^2+2^3]+3(2^4)\]

OpenStudy (freckles):

\[2S-S=(-1)(2)[1+2+2^2]+3(2^4) \\ 2S-S=(-1)(2)[2^{2}-1]+3(2^4)\\2S-S=(-1)(2)(2^2-1)+3(2^4) \\\]

OpenStudy (mathmath333):

ok i got that

OpenStudy (freckles):

\[S=-2(2^3)+2+3(2^4)=-2^4+3(2^4)+2=2\cdot 2^4+2\]

OpenStudy (freckles):

should be (2^3-1) above not (2^2-1)

OpenStudy (mathmath333):

ok thnks very much again

OpenStudy (freckles):

well I don't know how much help i was actually lol

OpenStudy (mathmath333):

lol me too